# Nonlinear analysis of causality for heat flow in heavy-ion collisions: constraints from equation of state

### v1

@@ I Introduction
The study of relativistic dissipative fluid dynamics has gained significant attention due to its relevance in describing extreme physical conditions such as those in the neutron stars, and ultra relativistic heavy-ion collisions. Among the most established frameworks for relativistic dissipation the oldest and widely used is the Mueller-Israel-Stewart (MIS) theory [1, 2], which addresses the acausality and instability issues inherent in the straightforward relativistic generalization of the Navier-Stokes equations [3, 4]. Among some of the recent improved formulation are DNMR theory [5], developed by Denicol, Niemi, Molnar, and Rischke; BRSSS theory developed by Baier, Romatschke, Son, Starinets, and Stephanov [6], and the BDNK framework, introduced by Bemfica, Disconzi, Noronha, and Kovtun [7].
@@
MIS theory, rooted in extended irreversible thermodynamics [10], introduces relaxation times for dissipative currents, ensuring a causal and stable evolution under specific conditions [6] and is one of the most widely used formulation of dissipative hydrodynamics in heavy-ion collisions.
@@
The importance of causality in relativistic hydrodynamics originates from the requirement that signal propagation must not exceed the speed of light. This condition is mathematically reflected in the hyperbolicity of the equations of motion, which is guaranteed when the eigenvalues of the characteristic determinant are real, positive and distinct [11]. Although many previous studies [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] have analyzed causality and stability mostly in the linear regime, including the effects of shear stress and net charge diffusion, there remains significant scope to explore the hyperbolicity of MIS theory in the context of heat conducting fluids, particularly in the non-linear regime.
@@
This issue is particularly relevant for systems exhibiting significant heat flux [29]. As noted by Hiscock and Lindblom [30], the strict hyperbolicity of the equations of motion can break down when the heat flux becomes extremely large, with the ratio $|q|/\varepsilon$ reaching some critical values. For instance, they reported that for $|q|/\varepsilon\approx 0.08898$ for a conformal fluid, the fluid equations cease to be hyperbolic, corresponding to a regime of extraordinarily large heat flux unlikely to be encountered in realistic physical systems. Additionally, they showed that varying transport coeffcients, such as adopting a ten times larger relaxation time through parameter $\beta_{1}$, can extend the hyperbolicity domain to $|q|/\varepsilon\approx 1/3$.
@@
In ultrarelativistic heavy-ion collisions, where deviations from equilibrium can be substantial, it is important to explore whether the heat flux values and causality limits predicted by the MIS framework fall within physically realistic ranges. One of the aims of this work is to estimate the typical order of magnitude of heat flux in such systems and investigate the extent of the causality-constrained parameter space. We show that with the current estimate of coefficient of thermal conductivity the heat flow could reach extremely large values for central fireball temperatures as low as 150 MeV. It is well known that the choice of hydrodynamic frame for defining the fluid four-velocity in dissipative fluids can lead to specific cases such as the Landau frame, where heat flow does not contribute to the energy flux $T^{0i}$ but reappears as a finite baryon current that is proportional to the thermal conductivity and the gradient of the ratio of chemical potential to temperature [31]. Conversely, one can choose a different frame, such as the Eckart frame, where the dissipative part of the baryon four-current vanishes and the energy flux becomes non-vanishing  [32, 33, 34].
@@ II Formulation
Where $\varepsilon$ is the energy density, $p$ is the thermodynamic pressure, related to $\varepsilon$ via an equation of state (EoS), and $\Pi$ is the bulk viscous pressure, representing the isotropic deviation of the pressure from its equilibrium value. The shear stress tensor $\pi^{\mu\nu}$ accounts for anisotropic deviations of the stress tensor and satisfies $\pi^{\mu\nu}u_{\nu}=0$ and $\pi^{\mu\nu}\Delta_{\mu\nu}=0$. The energy flux four-vector $w^{\mu}$ represents the flow of energy relative to the fluid’s rest frame and satisfies $w^{\mu}u_{\mu}=0$. In the conserved current $N^{\mu}$, $n$ is the density of the conserved quantity (e.g., net baryon), while $V^{\mu}$ is the particle diffusion current, representing deviations of the particle flux from the equilibrium rest frame, and it satisfies $V^{\mu}u_{\mu}=0$. The definitions of these quantities in terms of energy momentum tesnsor are as follows:
@@
We choose Eckart frame (where particle three current $V^{\mu}$ vanishes) for defining our hydrodynamic four velocity; by definition the particle diffusion $V^{\mu}$ is zero in this frame, and the energy flow $w^{\mu}$ coincides with the heat flow $q^{\mu}$. It is important to note that the coefficient of thermal conductivity $\kappa$ (defined below) may diverge for a QCD medium containing equal number of quark and antiquarks. In the limit where the quark chemical potential $\mu\ll T$, this divergence scales as $\sim 1/\mu^{2}$. Despite the divergent $\kappa$, the correction to the baryon current remains finite. This makes the Landau-Lifshitz frame a more suitable choice than the Eckart frame for systems produced at vanishingly small baryon chemical potential, as discussed in Ref. [31]. For simplicity, we neglect shear and bulk viscosity in this study. The conservation equations and the Muller-Israel-Stewart (MIS) evolution equation for heat flow reduce to the following form:
@@
For the one dimensional flow Eq.(1), and EoS $p=c_{s}^{2}\varepsilon=nT$ the conservation equatons Eq.(5) - Eq.(8) take the following form,
$\cosh\rho\frac{\partial n}{\partial t}+\sinh\rho\frac{\partial n}{\partial x}+n\sinh\rho\frac{\partial\rho}{\partial t}+n\cosh\rho\frac{\partial\rho}{\partial x}=0.$ (9)
@@
Figure 1: Characteristic velocities $v$ obtained from the numerical solution of Eq. (16) shown as a function of $q/\varepsilon$, for $\beta_{1}=\frac{5}{4p}$ and four different $c_{s}^{2}$=0.05, 0.10, 0.33, and 0.5.
III result
Figure 2: Same as Fig.(1) but for $\lambda$= 0.75.
We begin by fixing $\lambda=1$, which corresponds to $\beta_{1}=\frac{5}{4p}$, i.e., the second-order transport coefficient is set to its ultra-relativistic limit according to kinetic theory. The numerical roots for the characteristic velocities $v$ are computed for four different values of the squared speed of sound, $c_{s}^{2}=0.05,0.1,0.33,0.5$. These results are displayed in four panels of Fig. 1. Although a relativistic massless gas has $c_{s}^{2}=1/3$, one of the values choosen here $c_{s}^{2}=0.5$ is higher than this limit, we choose this value based on the fact that for extremely dense systems $c_{s}^{2}$ known to exceed the relativistic massless gas limit and may approach $c_{s}^{2}\rightarrow 1$ (see e.g. Kapusta and Gale [35], Weinberg[36], Bedaque and Steiner 2015 [37]).
From the top and bottom panels of Fig. 1, we observe that for all four scenarios, a causal region with four distinct real roots exists for $|\frac{q}{\varepsilon}|$ within a range $\approx$ 0.2 around the origin. In particular, the causal region expands to higher values of $|\frac{q}{\varepsilon}|$ as $c_{s}^{2}$ increases. This behavior clearly shows the dependence of causality on the equation of state through $c_{s}^{2}$.
The result for a massless ultra-relativistic gas ($c_{s}^{2}=\frac{1}{3}$) is shown in the bottom-left plot of Fig. 1. We observe that for $|q/\varepsilon|<0.18$, the characteristic speeds lie approximately in the range $0.45<v<0.63$. This finding is consistent with the results reported by Hiscock and Lindblom [30].
To further investigate the dependence of causality on the relaxation parameter $\lambda$, we consider two arbitrarily chosen values: a smaller $\lambda=0.75$ and a larger $\lambda=10$. We first examine the case $\lambda=0.75$, with the corresponding characteristic velocities shown in Fig. 2. Similar to the $\lambda=1$ case, we find that the causal region expands with increasing $c_{s}^{2}$. However, the range of $|\frac{q}{\varepsilon}|$ over which causality is preserved is somewhat narrower than in the $\lambda=1$ scenario, with a pronounced reduction in the causal region at smaller values of $c_{s}^{2}$. As we will see later, Lattice QCD calculations indicate that $c_{s}^{2}$ in the temperature range relevant for heavy-ion collisions reaches values around $\sim 0.15$ near the crossover temperature (see Fig. 4). This observation shows that the relaxation time, encoded in the parameter $\lambda$, also plays an important role in determining the system’s causality.
We now consider the case where $\lambda=10$. This value represents a significantly larger relaxation time compared to the previously analyzed cases ($\lambda=1$ and $\lambda=0.75$). The characteristic velocities $v$ as functions of $\frac{q}{\varepsilon}$ for four different values of $c_{s}^{2}=0.05,0.1,0.33$ , and $0.5$ are shown in Fig. 3.
From Fig. 3, we observe that the causal region becomes less constrained as $\lambda$ increases. Like other $\lambda$ values here also we observe for small values of $c_{s}^{2}$ (e.g., $c_{s}^{2}=0.05$ and $c_{s}^{2}=0.1$), the range of $|\frac{q}{\varepsilon}|$ for which all four roots remain real and distinct is significantly reduced compared to the cases of larger $c_{s}^{2}$. This indicates that although a larger relaxation time tends to instate causality in the system, the equation of state imposes constraint making it non-trivial to apply second-order hydrodynamics for systems with large gradients.
For larger values of $c_{s}^{2}$, such as $c_{s}^{2}=0.33$ (the ultra-relativistic limit) and $c_{s}^{2}=0.5$, the causal region remains relatively robust, and allows a very large range of $\frac{q}{\varepsilon}$ compared to smaller cases $\lambda$. Since MIS theory consider the dissipative corrections are to be smaller, it is not known exactly how large the heat flow could be considered before the theory breaks down. Notably, for $c_{s}^{2}=0.33$, the characteristic velocities lie within the sub-luminal range ($v<1$) only for $|q/\varepsilon|<0.25$. This suggests that in the ultra-relativistic regime, larger relaxation time ensure causality in systems with a large gradient of temperature.
@@
Figure 3: Characteristic velocities $v$ as functions of $\frac{q}{\varepsilon}$ for $\lambda=10$ and four different values of $c_{s}^{2}$: $0.05,0.1,0.33,0.5$. The causal regions are showing a significant enhancement in the range of $\frac{q}{\varepsilon}$ compared to smaller $\lambda$ cases.
Our findings shows the critical role of the relaxation parameter $\lambda$ as well as EoS in determining the causality (and hence the stability) of relativistic fluid with non-zero heat flow. While moderate values of $\lambda$ (e.g., $\lambda=1$) allow for broader causal regions, smaller values of $\lambda$ severely restrict the range of $\frac{q}{\varepsilon}$ for which causality is preserved. This behavior is demanding the importance of carefully calibrating $\lambda$ based on the physical properties of the system under study.
The results presented above demonstrate the important interplay between the equation of state through $c_{s}^{2}$, the relaxation parameter $\lambda$, and the relative magnitude of the heat flow $\frac{q}{\varepsilon}$ in determining the causal structure of relativistic fluids. Our findings confirm that causality is preserved only within specific ranges of $\frac{q}{\varepsilon}$, which depend sensitively on both $c_{s}^{2}$ and $\lambda$.
For $c_{s}^{2}=\frac{1}{3}$, which corresponds to the ultra-relativistic limit, the characteristic velocities remain subluminal ($v<1$) for small values of $|q/\varepsilon|$. The expansion of the causal region with increasing $c_{s}^{2}$ suggests that stiffer equations of state are more conducive to maintaining causality.
@@ III.1 $c_{s}^{2}(\varepsilon)$ from Lattice QCD result
Figure 5: Characteristic velocities $v$ as functions of $\frac{q}{\varepsilon}$ for three different $\lambda=$ 0.5, 1.0, and 20 and for speed of sound obtained from lattice QCD equation of state. The causal regions are showing a significant changes compared to the temperature independent constant speed of sound case.

### v2

@@ I Introduction
The study of relativistic dissipative fluid dynamics has gained significant attention due to its relevance in describing extreme physical conditions such as those in neutron stars and ultra-relativistic heavy-ion collisions. Among the most established frameworks for relativistic dissipation, the oldest and most widely used is the Mueller-Israel-Stewart (MIS) theory [1, 2], which addresses the acausality and instability issues inherent in the straightforward relativistic generalization of the Navier-Stokes equations [3, 4]. Among some of the recent improved formulations are the DNMR theory [5], developed by Denicol, Niemi, Molnar, and Rischke; the BRSSS theory developed by Baier, Romatschke, Son, Starinets, and Stephanov [6]; and the BDNK framework, introduced by Bemfica, Disconzi, Noronha, and Kovtun [7].
@@
MIS theory, rooted in extended irreversible thermodynamics [10], introduces relaxation times for dissipative currents, ensuring a causal and stable evolution under specific conditions [6] and is one of the most widely used formulations of dissipative hydrodynamics in heavy-ion collisions.
@@
The importance of causality in relativistic hydrodynamics originates from the requirement that signal propagation must not exceed the speed of light. This condition is mathematically reflected in the hyperbolicity of the equations of motion, which is guaranteed when the eigenvalues of the characteristic determinant are real, positive and distinct [11]. Although many previous studies [12, 13, 14, 15, 16, 17, 18, 19, 29, 20, 21, 22, 23, 24, 25, 26, 27, 28] have analyzed causality and stability mostly in the linear regime, including the effects of shear stress and net charge diffusion, there remains significant scope to explore the hyperbolicity and causality of MIS theory in the context of heat conducting fluids, particularly in the non-linear regime. Recent extensions have begun addressing nonlinear causality in contexts including diffusion, magnetohydrodynamics, and constraints on initial conditions for one-dimensional expanding fluids [30, 31, 28].
@@
This issue is particularly relevant for systems exhibiting significant heat flux [32]. As noted by Hiscock and Lindblom [33], the strict hyperbolicity of the equations of motion can break down when the heat flux becomes extremely large, with the ratio $|q|/\varepsilon$ reaching some critical values. For instance, they reported that for $|q|/\varepsilon\approx 0.08898$ for a conformal fluid, the fluid equations cease to be hyperbolic, corresponding to a regime of extraordinarily large heat flux unlikely to be encountered in realistic physical systems. Additionally, they showed that varying transport coefficients, such as adopting a ten times larger relaxation time through parameter $\beta_{1}$, can extend the hyperbolicity domain to $|q|/\varepsilon\approx 1/3$.
@@
In ultrarelativistic heavy-ion collisions, where deviations from equilibrium can be substantial, it is important to explore whether the heat flux values and causality limits predicted by the MIS framework fall within physically realistic ranges. One of the aims of this work is to estimate the typical order of magnitude of heat flux in such systems and investigate the extent of the causality-constrained parameter space. We show that with the current estimate of the coefficient of thermal conductivity the heat flow could reach extremely large values for central fireball temperatures as low as 150 MeV. It is well known that the choice of hydrodynamic frame for defining the fluid four-velocity in dissipative fluids can lead to specific cases such as the Landau frame, where heat flow does not contribute to the energy flux $T^{0i}$ but reappears as a finite baryon current that is proportional to the thermal conductivity and the gradient of the ratio of chemical potential to temperature [34]. Conversely, one can choose a different frame, such as the Eckart frame, where the dissipative part of the baryon four-current vanishes and the energy flux becomes non-vanishing [4, 35, 36].
@@ II Formulation
Here, $\varepsilon$ is the energy density, $p$ is the thermodynamic pressure, related to $\varepsilon$ via an equation of state (EoS), and $\Pi$ is the bulk viscous pressure, representing the isotropic deviation of the pressure from its equilibrium value. The shear stress tensor $\pi^{\mu\nu}$ accounts for anisotropic deviations of the stress tensor and satisfies $\pi^{\mu\nu}u_{\nu}=0$ and $\pi^{\mu\nu}\Delta_{\mu\nu}=0$. The energy flux four-vector $w^{\mu}$ represents the flow of energy relative to the fluid’s rest frame and satisfies $w^{\mu}u_{\mu}=0$. In the conserved current $N^{\mu}$, $n$ is the density of the conserved quantity (e.g., net baryon), while $V^{\mu}$ is the particle diffusion current, representing deviations of the particle flux from the equilibrium rest frame, and it satisfies $V^{\mu}u_{\mu}=0$. The definitions of these quantities in terms of energy-momentum tensor are as follows:
@@
We choose Eckart frame (where particle three current $V^{\mu}$ vanishes) for defining our hydrodynamic four-velocity; by definition the particle diffusion $V^{\mu}$ is zero in this frame, and the energy flow $w^{\mu}$ coincides with the heat flow $q^{\mu}$. It is important to note that the coefficient of thermal conductivity $\kappa$ (defined below) may diverge for a QCD medium containing equal number of quark and antiquarks. In the limit where the quark chemical potential $\mu\ll T$, this divergence scales as $\sim 1/\mu^{2}$. Despite the divergent $\kappa$, the correction to the baryon current remains finite. This makes the Landau-Lifshitz frame a more suitable choice than the Eckart frame for systems produced at vanishingly small baryon chemical potential, as discussed in Ref. [34]. For simplicity, we neglect shear and bulk viscosity in this study. The conservation equations and the Muller-Israel-Stewart (MIS) evolution equation for heat flow reduce to the following form:
@@
For the one dimensional flow Eq.(1), and EoS $p=c_{s}^{2}\varepsilon=nT$ the conservation equations Eq.(5) - Eq.(8) take the following form,
$\cosh\rho\frac{\partial n}{\partial t}+\sinh\rho\frac{\partial n}{\partial x}+n\sinh\rho\frac{\partial\rho}{\partial t}+n\cosh\rho\frac{\partial\rho}{\partial x}=0,$ (9)
@@
Figure 1: Characteristic velocities $v$ obtained from the numerical solution of Eq. (16) shown as a function of $q/\varepsilon$, for $\beta_{1}=\frac{5}{4p}$ and four different $c_{s}^{2}$=0.05, 0.10, 0.33, and 0.5. The shaded green region indicates the causal hyperbolic region.
III Results
We analyze the causal structure of the MIS theory by numerically solving the characteristic equation, Eq. (16), for the propagation velocities $v$. The system is hyperbolic and causal when all four roots are real and subluminal ($|v|\leq 1$). We investigate the causal parameter space as a function of the relaxation time parameter, $\lambda$ (where $\beta_{1}=\lambda\frac{5}{4p}$), the equation of state (EoS), and the normalized heat flux, $q/\varepsilon$.
For the EoS, we consider several constant values for the squared speed of sound, $c_{s}^{2}$. While a relativistic massless gas corresponds to $c_{s}^{2}=1/3$, we also analyze stiffer cases up to $c_{s}^{2}=0.5$, as such high values may be reached in extremely dense systems [37, 38, 39].
Figures 1 and 2 show the characteristic velocities for $\lambda=1$ and $\lambda=10$. Two key trends are immediately apparent. First, for a fixed $\lambda$, a stiffer EoS (larger $c_{s}^{2}$) consistently expands the range of $|q/\varepsilon|$ for which causality is maintained. Second, comparing the two figures shows that a larger relaxation time (increasing $\lambda$ from 1 to 10) significantly enlarges the causal domain. For instance, in the ultra-relativistic limit ($c_{s}^{2}=1/3$), the causal region is limited to $|q/\varepsilon|\lesssim 0.18$ for $\lambda=1$ (consistent with Hiscock and Lindblom [33]), but expands to $|q/\varepsilon|\lesssim 0.25$ for $\lambda=10$.
To visualize these dependencies more globally, Fig. 3 maps the causality regions in the ($q/\varepsilon$, $c_{s}^{2}$) plane for $\lambda=0.75,1.0,$ and $10.0$. These plots confirm that the causal domain (green shaded area) is highly sensitive to both the EoS and the relaxation time. The region of causal evolution shrinks dramatically for softer equations of state and smaller values of $\lambda$. This strong interplay shows the important role of the second-order transport coefficients and the EoS in maintaining a well-posed hydrodynamic theory, particularly in regimes with large temperature gradients where the heat flux may be significant.
@@ III Results
Figure 2: Characteristic velocities $v$ as functions of $\frac{q}{\varepsilon}$ for $\lambda=10$ and four different values of $c_{s}^{2}$: $0.05,0.1,0.33,0.5$. The causal regions, marked by the shaded green area, show a significant enhancement in the range of $\frac{q}{\varepsilon}$ compared to smaller $\lambda$ cases.
Figure 3: Causality regions in the ($q/\varepsilon$, $c_{s}^{2}$) plane for different values of the relaxation parameter, $\lambda=0.75,1.0,$ and $10.0$. The green (Hyperbolic/ Causal), yellow (Hyperbolic/Acausal), and red (Non-hyperbolic/Acausal) regions denote the parameter space where the characteristic velocities are all real and subluminal, all real but at least one is superluminal, and not all real, respectively.
@@ III.1 $c_{s}^{2}(\varepsilon)$ from Lattice QCD result
Figure 5: Characteristic velocities $v$ as functions of $q/\varepsilon$ for three different values of $\lambda=0.5,1.0,$ and $20.0$, calculated using a lattice QCD equation of state. The plots show the results for positive $q/\varepsilon$, with the causal hyperbolic region highlighted by green rectangles.